A student hears 6 beat in 2.0 s two different notes.If one note has a frequency of 430 hz what are the possible frequencies of second note?


1
Expert's answer
2020-06-29T14:04:13-0400
"\Delta f=\frac{6}{2}=3\ Hz"

"f_1=430-3=427\ Hz,\\f_2=430+3=433\ Hz"